OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)?

OpenStudy (anonymous):

Use (y2 - y1) / (x2 - x1) = m and then put the equation into the point-slope formula of y = mx + b.

OpenStudy (anonymous):

(-3,-9) right?

OpenStudy (anonymous):

then i put it in point slope formula

OpenStudy (anonymous):

Where di you get (-3,-9) ? You have to use the fraction I gave you in the first post.

OpenStudy (anonymous):

so then the equation would be written out 10-1/1-(-2)?

OpenStudy (anonymous):

Yes, that will = the slope or m. Then you can take one of the points and put it into y = mx + b with your numerical value for m. The only variable left will be b. Solve for b. Then, you can go back to y = mx + b with the values of m and b and you will have your equation.

OpenStudy (anonymous):

point-slope form is y-y1=M(x-x1)...